OpenStudy (anonymous):

Solve each system of equations by graphing 2x+5y=5 y=2/5x+5

OpenStudy (anonymous):

Isolate the variable y to the left-hand side of the equation for the first equation. Then you have two equations with y isolated and as functions of x. Then you can graph them.

OpenStudy (amtran_bus):

If you have a Ti 84 or Nspire, you can graph them on it and the answer will be the intercept(s).

OpenStudy (anonymous):

Nope i don't have either

OpenStudy (amtran_bus):

Wait, plug in the y value your given into the first equation.

OpenStudy (amtran_bus):

Have you heard of wolfram alpha?

OpenStudy (amtran_bus):

@overtonka , yes I tagged you lol

OpenStudy (amtran_bus):

What calc do you have?

OpenStudy (anonymous):

i don't have one

OpenStudy (amtran_bus):

There is your (x,y) coord.

OpenStudy (amtran_bus):

@overtonka you can use that site as a calc.

OpenStudy (anonymous):

thanks

OpenStudy (amtran_bus):

Of course, aero was right also. You could change it to y=mx+b form and graph them on paper. Do you really understand?

OpenStudy (anonymous):

no lol

OpenStudy (amtran_bus):

I want to be a teach, so ok. Here we go.

OpenStudy (amtran_bus):

What method do you want to use?

OpenStudy (anonymous):

i think that i am best at elimination

OpenStudy (amtran_bus):

Can you tell me the first step in that?

OpenStudy (anonymous):

get the same numbers

OpenStudy (amtran_bus):

Give me a sec...

OpenStudy (amtran_bus):

@Aeronautix can you take over from here, I need to go and dont want to leave her hanging.

jimthompson5910 (jim_thompson5910):

Using substitution, we get.. 2x+5y=5 2x+5(2/5x + 5)=5 ... plug in y = 2/5x+5 ... ie replace y with 2/5x+5 2x + 2x + 25 = 5 ... distribute 4x + 25 = 5 4x = 5 - 25 4x = -20 x = -20/4 x = -5 So the x part of the solution is x = -5. You would then use this x value to find y y = (2/5)*(-5) + 5 y = -10/5 + 5 y = -2 + 5 y = 3 and the y part of the solution is y = 3 The graph will confirm this visually because the two lines intersect at (-5,3)

OpenStudy (anonymous):

A plot is attached.